# Consensus over Clustered Networks Using Intermittent and Asynchronous Output Feedback

### v2

@@ IV Hybrid System Development
$\dot{x}_{p}$ $=Ax_{p}+BK_{u}H\sum_{q\in\mathcal{N}_{p}}(x_{q}-x_{p})+BK_{u}\sum_{r\in[M^{\ast}]}\tilde{\zeta}_{pr}$ (13)
$\phantom{=}\ +BK_{u}\tilde{\eta}_{p},$
@@
$\dot{x}$ $=((I_{N}\otimes A)-(\mathsf{L}\otimes BK_{u}H))x+(I_{N}\otimes BK_{u})\tilde{\eta}$ (14)
$\phantom{=}\ +(I_{N}\otimes BK_{u})(1_{M^{\ast}}^{\top}\otimes I_{mN})\tilde{\zeta}.$
Recall the objective, that is, drive $x$ from any initial condition to a configuration where $x_{p}=x_{q}$ for all distinct $p,q\in\mathcal{V}$. Since any configuration $x$ can be decomposed into orthogonal components, where $x=x^{\|}+x^{\perp}$, $x^{\|}\triangleq((1_{N}1_{N}^{\top})/N\otimes I_{n})x$ is the part of $x$ in consensus, and $x^{\perp}\triangleq(\mathsf{S}\otimes I_{n})x$ is the part of $x$ in dissensus, one can achieve the objective by driving $\|x^{\perp}\|$ to 00. Naturally, we attempted to produce a GES result using the dynamics of $x^{\perp}$. However, in this particular setting, such an exercise was unfruitful because of the interaction between the non-trivial kernel of $\mathsf{L}$ and the linear time-invariant dynamics in (5).
The use of $\mathsf{S}\mathsf{L}=\mathsf{L}\mathsf{S}$ and the substitution of (14) into the time derivative of $x^{\perp}$ yields a linear ordinary differential equation (ODE) that depends on the variable $(x^{\perp},\tilde{\eta},\tilde{\zeta})$. The matrix that pre-multiplies $x^{\perp}$ cannot, in general, be made into a stability matrix through the selection of $K_{u}$, unless $A$ itself is a stability matrix. Fortunately, Lemma 1 enables the succeeding result, which motivates the analysis of an alternative ODE.
@@
$\|z^{\perp}\|^{2}$ $=\|(\mathsf{S}\otimes I_{n})z\|^{2}=z^{\top}(\mathsf{S}\otimes I_{n})^{\top}(\mathsf{S}\otimes I_{n})z$
$=z^{\top}(\mathsf{S}^{\top}\otimes I_{n})(\mathsf{S}\otimes I_{n})z=z^{\top}(\mathsf{S}^{\top}\mathsf{S}\otimes I_{n})z$
$=z^{\top}(\mathsf{S}^{2}\otimes I_{n})z=z^{\top}(\mathsf{S}\otimes I_{n})z$
$=z^{\top}(\mathsf{V}\mathsf{V}^{\top}\otimes I_{n})z=z^{\top}(\mathsf{V}\otimes I_{n})(\mathsf{V}^{\top}\otimes I_{n})z$
$=\|(\mathsf{V}^{\top}\otimes I_{n})z\|^{2},$
@@
Let $x^{\circ}\triangleq(\mathsf{V}^{\top}\otimes I_{n})x\in\mathbb{R}^{n(N-1)}$, and recall that consensus in $\{x_{p}\}_{p\in\mathcal{V}}$ is achieved whenever a solution of (14) converges to the agreement subspace defined by
$\{x\in\mathbb{R}^{nN}\colon\forall_{p,q\in\mathcal{V}}\ x_{p}=x_{q}\}.$ (15)
Since the distance of an arbitrary configuration $x$ to the agreement subspace in (15) is quantified by $\|x^{\perp}\|$ and $\|x^{\perp}\|=\|x^{\circ}\|$ by Lemma 2, $\|x^{\circ}\|$ is an alternative metric for consensus. Therefore, $x_{p}=x_{q}$ for all distinct $p,q\in\mathcal{V}$ if and only if $\|x^{\circ}\|=0$.
@@
$\dot{x}^{\circ}$ $=((I_{N-1}\otimes A)-(\mathsf{D}\otimes BK_{u}H))x^{\circ}+(\mathsf{V}^{\top}\otimes BK_{u})\tilde{\eta}$ (16)
$\phantom{=}\ +(\mathsf{V}^{\top}\otimes BK_{u})(1_{M^{\ast}}^{\top}\otimes I_{mN})\tilde{\zeta}.$
Observe the matrix $(I_{N-1}\otimes A)-(\mathsf{D}\otimes BK_{u}H)$ can be made stable through the selection of $K_{u}$, independent of the stability of $A$, provided $(A,B)$ is controllable, $(A,H)$ is observable, and $K_{u}H\neq 0_{d\times n}$ given [37, Lemma 2.2 & Theorem 2.6]. [^6] Let $\lambda_{1}>0$ denote the smallest eigenvalue of $\mathsf{D}$. Using a partial order on the space of positive semi definite matrices, one can show that $(I_{N-1}\otimes A)-(\mathsf{D}\otimes BK_{u}H)$ is a stability matrix provided $A-\lambda_{1}BK_{u}H$ is a stability matrix. Such a matrix $K_{u}$ can be computed using[37, Theorem 2.6]. Consider the block diagonal matrix $\mathtt{e}\triangleq\text{diag}(\mathtt{e}_{1},\mathtt{e}_{2},...,\mathtt{e}_{N})\in\mathbb{R}^{N^{2}\times N}$, where $\mathtt{e}_{k}$ denotes the $k^{\text{th}}$ standard basis vector in $\mathbb{R}^{N}$. Substituting $\mathsf{L}=\mathsf{L}\mathsf{S}$, (2), the output equation in (5), (11), and $x^{\circ}\triangleq(\mathsf{V}^{\top}\otimes I_{n})x$ into $\tilde{\eta}$ for every $p\in\mathcal{V}$ yields
@@
where $\mathsf{L}_{0}\triangleq\text{diag}(\mathsf{L}[\mathcal{C}_{1}],\mathsf{L}[\mathcal{C}_{2}],...,\mathsf{L}[\mathcal{C}_{N}])\in\mathbb{R}^{N^{2}\times N^{2}}$ and $\mathsf{C}\triangleq(I_{N}\otimes H)(\mathtt{e}^{\top}\mathsf{L}_{0}(I_{N}\otimes\mathsf{V})\otimes I_{n})(1_{N}\otimes I_{n(N-1)})\in\mathbb{R}^{mN\times n(N-1)}$. The substitution of the flow equation in (8) for each $p\in\mathcal{V}$, (16), and (17) into the time derivative of (17) yields
$\dot{\tilde{\eta}}$ $=(\mathsf{C}((I_{N-1}\otimes A)-(\mathsf{D}\otimes BK_{u}H))-K_{\eta}\mathsf{C})x^{\circ}$ (18)
$\phantom{=}\ +(\mathsf{C}(\mathsf{V}^{\top}\otimes BK_{u})+K_{\eta})\tilde{\eta}$
$\phantom{=}\ +\mathsf{C}(\mathsf{V}^{\top}\otimes BK_{u})(1_{M^{\ast}}^{\top}\otimes I_{mN})\tilde{\zeta}$
with $K_{\eta}\triangleq\text{diag}(K_{\eta,1},K_{\eta,2},...,K_{\eta,N})\in\mathbb{R}^{mN\times mN}$. Further, the substitution of $\mathsf{L}=\mathsf{L}\mathsf{S}$, (2), the output equation in (5), (12), and $x^{\circ}\triangleq(\mathsf{V}^{\top}\otimes I_{n})x$ into $\tilde{\zeta}$ for every $p\in\mathcal{V}$ and $r\in[M^{\ast}]$ yields
@@
where $\mathsf{I}_{r}\triangleq(I_{N}\otimes H)(\mathtt{e}^{\top}(I_{N}\otimes\mathsf{L}[\mathcal{V}^{r}]\mathsf{V})\otimes I_{n})(1_{N}\otimes I_{n(N-1)})\in\mathbb{R}^{mN\times n(N-1)}$. Substituting (19) for every $r\in[M^{\ast}]$ into $\tilde{\zeta}$ yields
@@
where $\mathsf{I}\triangleq\text{diag}(\mathsf{I}_{1},\mathsf{I}_{2},...,\mathsf{I}_{M^{\ast}})(1_{M^{\ast}}\otimes I_{n(N-1)})$. The substitution of the flow equation in (9) for every $p\in\mathcal{V}$ and $r\in[M^{\ast}]$, (16), and (20) into the time derivative of (20) yields
$\dot{\tilde{\zeta}}$ $=(\mathsf{I}((I_{N-1}\otimes A)-(\mathsf{D}\otimes BK_{u}H))-K_{\zeta}\mathsf{I})x^{\circ}$ (21)
$\phantom{=}\ +\mathsf{I}(\mathsf{V}^{\top}\otimes BK_{u})\tilde{\eta}$
$\phantom{=}\ +(\mathsf{I}(\mathsf{V}^{\top}\otimes BK_{u})(1_{M^{\ast}}^{\top}\otimes I_{mN})+K_{\zeta})\tilde{\zeta},$
where $K_{\zeta,r}\triangleq\text{diag}(K_{\zeta,1r},K_{\zeta,2r},...,K_{\zeta,Nr})\in\mathbb{R}^{mN\times mN}$ and $K_{\zeta}\triangleq\text{diag}(K_{\zeta,1},K_{\zeta,2},...,K_{\zeta,M^{\ast}})\in\mathbb{R}^{mNM^{\ast}\times mNM^{\ast}}$.
Let $z\triangleq(x^{\circ},\tilde{\eta},\tilde{\zeta})\in\mathbb{R}^{n(N-1)+mN+mNM^{\ast}}$. By leveraging the ODEs in (16), (18), and (21), a dynamical system can be developed for the variable $z$ during flows, namely,
@@
$\mathsf{F}$ $\triangleq\left[\begin{array}[]{ccc}\mathsf{F}_{11}&\mathsf{F}_{12}&\mathsf{F}_{13}\\ \mathsf{C}\mathsf{F}_{11}-K_{\eta}\mathsf{C}&\mathsf{C}\mathsf{F}_{12}+K_{\eta}&\mathsf{C}\mathsf{F}_{13}\\ \mathsf{I}\mathsf{F}_{11}-K_{\zeta}\mathsf{I}&\mathsf{I}\mathsf{F}_{12}&\mathsf{I}\mathsf{F}_{13}+K_{\zeta}\end{array}\right],$
$\mathsf{F}_{11}$ $\triangleq(I_{N-1}\otimes A)-(\mathsf{D}\otimes BK_{u}H)$
$\mathsf{F}_{12}$ $\triangleq\mathsf{V}^{\top}\otimes BK_{u},\quad\mathsf{F}_{13}\triangleq(\mathsf{V}^{\top}\otimes BK_{u})(1_{M^{\ast}}^{\top}\otimes I_{mN}).$
Let $\mathcal{H}$ denote the closed-loop hybrid system for the ensemble. The state variable and state space of $\mathcal{H}$ are denoted by $\xi\triangleq(z,\tau,\rho)\in\mathcal{X}$ and $\mathcal{X}\triangleq\mathbb{R}^{n(N-1)}\times\mathbb{R}^{mN}\times\mathbb{R}^{mNM^{\ast}}\times\mathcal{T}\times\mathcal{R}$, respectively. Observe $\mathcal{T}\triangleq[0,T_{2}^{1}]\times[0,T_{2}^{2}]\times...\times[0,T_{2}^{N}]$ and $\mathcal{R}\triangleq[0,T_{4}^{1}]\times[0,T_{4}^{2}]\times...\times[0,T_{2}^{M^{\ast}}]$. For every $p\in\mathcal{V}$ and $r\in[M^{\ast}]$, let
$C_{p}\triangleq\{\xi\in\mathcal{X}\colon\tau_{p}\in[0,T_{2}^{p}]\},\ C^{r}\triangleq\{\xi\in\mathcal{X}\colon\rho_{r}\in[0,T_{4}^{r}]\}.$
The flow set of $\mathcal{H}$ is
$C\triangleq\big{(}\cap_{p\in\mathcal{V}}C_{p}\big{)}\bigcap\big{(}\cap_{r\in[M^{\ast}]}C^{r}\big{)}.$ (23)
@@
$f(\xi)\triangleq\begin{bmatrix}\mathsf{F}z\\ -1_{N}\\ -1_{M^{\ast}}\end{bmatrix}.$ (24)
The flow map in (24) follows from the substitution of the flow equation in (6) for each $p\in\mathcal{V}$, the flow equation in (7) for each $r\in[M^{\ast}]$, and the dynamics of $z$ in (22) into the time derivative of $\xi$. For each $p\in\mathcal{V}$ and $r\in[M^{\ast}]$, let
$D_{p}\triangleq\{\xi\in\mathcal{X}\colon\tau_{p}=0\},\ D^{r}\triangleq\{\xi\in\mathcal{X}\colon\rho_{r}=0\}.$
The jump set of $\mathcal{H}$ is
$D\triangleq\big{(}\cup_{p\in\mathcal{V}}D_{p}\big{)}\bigcup\big{(}\cup_{r\in[M^{\ast}]}D^{r}\big{)}.$ (25)
The differential inclusion governing the jumps of $\mathcal{H}$ is $\xi^{+}\in G(\xi)$, where the set-valued jump map $G\colon\mathcal{X}\rightrightarrows\mathcal{X}$ is
$G(\xi)$$\triangleq\{G_{p}(\xi)\colon\xi\in D_{p}\text{ for some }p\in\mathcal{V}\}$ (26)
$\phantom{\triangleq}\ \cup\{G^{r}(\xi)\colon\xi\in D^{r}\text{ for some }r\in[M^{\ast}]\},$
$G_{p}(\xi)$ $\triangleq\left[\begin{array}[]{c}x^{\circ}\\ \left[\tilde{\eta}_{1}^{\top},...,\tilde{\eta}_{p-1}^{\top},0_{m}^{\top},\tilde{\eta}_{p+1}^{\top},...,\tilde{\eta}_{N}^{\top}\right]^{\top}\\ \tilde{\zeta}\\ \left[\tau_{1},...,\tau_{p-1},[T_{1}^{p},T_{2}^{p}],\tau_{p+1},...,\tau_{N}\right]^{\top}\\ \rho\end{array}\right],$
$G^{r}(\xi)$ $\triangleq\left[\begin{array}[]{c}x^{\circ}\\ \tilde{\eta}\\ \left[\tilde{\zeta}_{1}^{\top},...,\tilde{\zeta}_{r-1}^{\top},0_{mN}^{\top},\tilde{\zeta}_{r+1}^{\top},...,\tilde{\zeta}_{M^{\ast}}^{\top}\right]^{\top}\\ \tau\\ \left[\rho_{1},...,\rho_{r-1},[T_{3}^{r},T_{4}^{r}],\rho_{r+1},...,\rho_{M^{\ast}}\right]^{\top}\end{array}\right].$
@@
$\mathcal{A}\triangleq\big{\{}\xi\in\mathcal{X}\colon(x^{\circ},\tilde{\eta},\tilde{\zeta})=(0_{n(N-1)},0_{mN},0_{mNM^{\ast}})\big{\}}$ (27)
@@
Under the construction of $\mathcal{H}$, the flow set $C$ and jump set $D$ are closed. Additionally, the flow map $f$ is continuous. The jump map $G$ is outer semi-continuous and locally bounded. As a result, $\mathcal{H}$ satisfies the hybrid basic conditions [32, Assumption 6.5], and [32, Theorem 6.8] implies $\mathcal{H}$ is nominally well-posed.
V Stability Analysis

### v3

@@ 4.3 Hybrid System Modeling
$\dot{x}_{p}=Ax_{p}+BK_{u}H\sum_{q\in\mathcal{N}_{p}}(x_{q}-x_{p})+BK_{u}\sum_{r\in[M^{\ast}]}\tilde{\zeta}_{pr}+BK_{u}\tilde{\eta}_{p},$ (15)
@@
$\dot{x}=((I_{N}\otimes A)-(\mathsf{L}\otimes BK_{u}H))x+(I_{N}\otimes BK_{u})\tilde{\eta}+(I_{N}\otimes BK_{u})(1_{M^{\ast}}^{\top}\otimes I_{mN})\tilde{\zeta}$. (16)
Recall the objective: drive $x$ from any initial condition to a configuration where $x_{p}=x_{q}$ for all distinct $p,q\in\mathcal{V}$. Since any configuration $x$ can be decomposed into orthogonal components, where $x=x^{\|}+x^{\perp}$, $x^{\|}\coloneqq((1_{N}1_{N}^{\top})/N\otimes I_{n})x$ is the part of $x$ in consensus, and $x^{\perp}\coloneqq(\mathsf{S}\otimes I_{n})x$ is the part of $x$ in dissensus, one can achieve the objective by driving $\|x^{\perp}\|$ to 0. Naturally, we attempted to produce a GES result using the dynamics of $x^{\perp}$. However, in this particular setting, such an exercise was unfruitful because of the interaction between the non-trivial kernel of the graph Laplacian matrix $\mathsf{L}$ and the linear time-invariant dynamics in (5).
[…]
@@
$\|z^{\perp}\|^{2}$ $=\|(\mathsf{S}\otimes I_{n})z\|^{2}=z^{\top}(\mathsf{S}\otimes I_{n})^{\top}(\mathsf{S}\otimes I_{n})z=z^{\top}(\mathsf{S}^{\top}\otimes I_{n})(\mathsf{S}\otimes I_{n})z=z^{\top}(\mathsf{S}^{\top}\mathsf{S}\otimes I_{n})z$
$=z^{\top}(\mathsf{S}^{2}\otimes I_{n})z=z^{\top}(\mathsf{S}\otimes I_{n})z=z^{\top}(\mathsf{V}\mathsf{V}^{\top}\otimes I_{n})z=z^{\top}(\mathsf{V}\otimes I_{n})(\mathsf{V}^{\top}\otimes I_{n})z=\|(\mathsf{V}^{\top}\otimes I_{n})z\|^{2},$
@@
Let $x^{\circ}\coloneqq(\mathsf{V}^{\top}\otimes I_{n})x\in\mathbb{R}^{n(N-1)}$, and recall that consensus in $\{x_{p}\}_{p\in\mathcal{V}}$ is achieved whenever a solution of (16) converges to the agreement subspace defined by $\mathtt{A}\coloneqq\{x\in\mathbb{R}^{nN}\colon\forall_{p,q\in\mathcal{V}}\ x_{p}=x_{q}\}$. Since the distance of an arbitrary configuration $x$ to the agreement subspace $\mathtt{A}$ is quantified by $\|x^{\perp}\|$ and $\|x^{\perp}\|=\|x^{\circ}\|$ by Lemma 2, $\|x^{\circ}\|$ is an alternative metric for consensus. Therefore, $x_{p}=x_{q}$ for all distinct $p,q\in\mathcal{V}$ if and only if $\|x^{\circ}\|=0$.
@@
$\dot{x}^{\circ}=((I_{N-1}\otimes A)-(\mathsf{D}\otimes BK_{u}H))x^{\circ}+(\mathsf{V}^{\top}\otimes BK_{u})\tilde{\eta}+(\mathsf{V}^{\top}\otimes BK_{u})(1_{M^{\ast}}^{\top}\otimes I_{mN})\tilde{\zeta}$. (17)
Observe, the matrix $(I_{N-1}\otimes A)-(\mathsf{D}\otimes BK_{u}H)$ can be made stable through the selection of $K_{u}$, independent of the stability of $A$, provided $(A,B)$ is controllable, $(A,H)$ is observable, and $K_{u}H\neq 0_{d\times n}$ given [Rodrigues2022, Lemma 2.2 & Theorem 2.6]. [^9] Let $\lambda_{1}>0$ denote the smallest eigenvalue of $\mathsf{D}$. Using a partial order on the space of positive semi definite matrices, one can show that $(I_{N-1}\otimes A)-(\mathsf{D}\otimes BK_{u}H)$ is a stability matrix provided $A-\lambda_{1}BK_{u}H$ is a stability matrix. Such a matrix $K_{u}$ can be computed using[Rodrigues2022, Theorem 2.6]. Consider the block diagonal matrix $\mathtt{e}\coloneqq\text{diag}(\mathtt{e}_{1},\mathtt{e}_{2},...,\mathtt{e}_{N})\in\mathbb{R}^{N^{2}\times N}$, where $\mathtt{e}_{k}$ denotes the $k^{\text{th}}$ standard basis vector in $\mathbb{R}^{N}$. Substituting $\mathsf{L}=\mathsf{L}\mathsf{S}$, (2), the output equation in (5), (13), and $x^{\circ}\coloneqq(\mathsf{V}^{\top}\otimes I_{n})x$ into $\tilde{\eta}$ for every $p\in\mathcal{V}$ yields
@@
where $\mathsf{L}_{0}\coloneqq\text{diag}(\mathsf{L}[\mathcal{C}_{1}],\mathsf{L}[\mathcal{C}_{2}],...,\mathsf{L}[\mathcal{C}_{N}])\in\mathbb{R}^{N^{2}\times N^{2}}$ and $\mathsf{C}\coloneqq(I_{N}\otimes H)(\mathtt{e}^{\top}\mathsf{L}_{0}(I_{N}\otimes\mathsf{V})\otimes I_{n})(1_{N}\otimes I_{n(N-1)})\in\mathbb{R}^{mN\times n(N-1)}$. The substitution of the flow equation in (10) for each $p\in\mathcal{V}$, (17), and (18) into the time derivative of (18) yields
$\dot{\tilde{\eta}}=(\mathsf{C}((I_{N-1}\otimes A)-(\mathsf{D}\otimes BK_{u}H))-K_{\eta}\mathsf{C})x^{\circ}+(\mathsf{C}(\mathsf{V}^{\top}\otimes BK_{u})+K_{\eta})\tilde{\eta}+\mathsf{C}(\mathsf{V}^{\top}\otimes BK_{u})(1_{M^{\ast}}^{\top}\otimes I_{mN})\tilde{\zeta}$ (19)
with $K_{\eta}\coloneqq\text{diag}(K_{\eta,1},K_{\eta,2},...,K_{\eta,N})\in\mathbb{R}^{mN\times mN}$. Further, the substitution of $\mathsf{L}=\mathsf{L}\mathsf{S}$, (2), the output equation in (5), (14), and $x^{\circ}\coloneqq(\mathsf{V}^{\top}\otimes I_{n})x$ into $\tilde{\zeta}$ for every $p\in\mathcal{V}$ and $r\in[M^{\ast}]$ yields
@@
where $\mathsf{I}_{r}\coloneqq(I_{N}\otimes H)(\mathtt{e}^{\top}(I_{N}\otimes\mathsf{L}[\mathcal{V}^{r}]\mathsf{V})\otimes I_{n})(1_{N}\otimes I_{n(N-1)})\in\mathbb{R}^{mN\times n(N-1)}$. Substituting (20) for every $r\in[M^{\ast}]$ into $\tilde{\zeta}$ yields
@@
where $\mathsf{I}\coloneqq\text{diag}(\mathsf{I}_{1},\mathsf{I}_{2},...,\mathsf{I}_{M^{\ast}})(1_{M^{\ast}}\otimes I_{n(N-1)})$. The substitution of the flow equation in (11) for every $p\in\mathcal{V}$ and $r\in[M^{\ast}]$, (17), and (21) into the time derivative of (21) yields
$\dot{\tilde{\zeta}}=(\mathsf{I}((I_{N-1}\otimes A)-(\mathsf{D}\otimes BK_{u}H))-K_{\zeta}\mathsf{I})x^{\circ}+\mathsf{I}(\mathsf{V}^{\top}\otimes BK_{u})\tilde{\eta}+(\mathsf{I}(\mathsf{V}^{\top}\otimes BK_{u})(1_{M^{\ast}}^{\top}\otimes I_{mN})+K_{\zeta})\tilde{\zeta}$, (22)
where $K_{\zeta,r}\coloneqq\text{diag}(K_{\zeta,1r},K_{\zeta,2r},...,K_{\zeta,Nr})\in\mathbb{R}^{mN\times mN}$ and $K_{\zeta}\coloneqq\text{diag}(K_{\zeta,1},K_{\zeta,2},...,K_{\zeta,M^{\ast}})\in\mathbb{R}^{mNM^{\ast}\times mNM^{\ast}}$.
Let $z\coloneqq(x^{\circ},\tilde{\eta},\tilde{\zeta})\in\mathbb{R}^{n(N-1)+mN+mNM^{\ast}}$. By leveraging the ODEs in (17), (19), and (22), a dynamical system can be developed for the variable $z$ during flows, namely,
@@
$\mathsf{F}\coloneqq\left[\begin{array}[]{ccc}\mathsf{F}_{11}&\mathsf{F}_{12}&\mathsf{F}_{13}\\ \mathsf{C}\mathsf{F}_{11}-K_{\eta}\mathsf{C}&\mathsf{C}\mathsf{F}_{12}+K_{\eta}&\mathsf{C}\mathsf{F}_{13}\\ \mathsf{I}\mathsf{F}_{11}-K_{\zeta}\mathsf{I}&\mathsf{I}\mathsf{F}_{12}&\mathsf{I}\mathsf{F}_{13}+K_{\zeta}\end{array}\right],\quad\quad\begin{aligned} \mathsf{F}_{11}&\coloneqq(I_{N-1}\otimes A)-(\mathsf{D}\otimes BK_{u}H),\\ \mathsf{F}_{12}&\coloneqq\mathsf{V}^{\top}\otimes BK_{u},\\ \mathsf{F}_{13}&\coloneqq(\mathsf{V}^{\top}\otimes BK_{u})(1_{M^{\ast}}^{\top}\otimes I_{mN}).\end{aligned}$
Let $\mathcal{H}$ denote the closed-loop hybrid system for the ensemble. The state variable and state space of $\mathcal{H}$ are denoted by $\xi\coloneqq(z,\tau,\rho)\in\mathcal{X}$ and $\mathcal{X}\coloneqq\mathbb{R}^{n(N-1)}\times\mathbb{R}^{mN}\times\mathbb{R}^{mNM^{\ast}}\times\mathcal{T}\times\mathcal{R}$, respectively. Observe, $\mathcal{T}\coloneqq[0,T_{2}^{1}]\times[0,T_{2}^{2}]\times...\times[0,T_{2}^{N}]$ and $\mathcal{R}\coloneqq[0,T_{4}^{1}]\times[0,T_{4}^{2}]\times...\times[0,T_{2}^{M^{\ast}}]$. For every $p\in\mathcal{V}$ and $r\in[M^{\ast}]$, let $C_{\tau,p}\coloneqq\{\xi\in\mathcal{X}\colon\tau_{p}\in[0,T_{2}^{p}]\}$ and $C_{\rho,r}\coloneqq\{\xi\in\mathcal{X}\colon\rho_{r}\in[0,T_{4}^{r}]\}$. The flow set of $\mathcal{H}$ is
$C\coloneqq\big(\cap_{p\in\mathcal{V}}C_{\tau,p}\big)\cap\big(\cap_{r\in[M^{\ast}]}C_{\rho,r}\big).$ (24)
@@
$f(\xi)\coloneqq(\mathsf{F}z,-1_{N},-1_{M^{\ast}}).$ (25)
The flow map in (25) follows from the substitution of the flow equation in (8) for each $p\in\mathcal{V}$, the flow equation in (9) for each $r\in[M^{\ast}]$, and the dynamics of $z$ in (23) into the time derivative of $\xi$. For each $p\in\mathcal{V}$ and $r\in[M^{\ast}]$, let $D_{\tau,p}\coloneqq\{\xi\in\mathcal{X}\colon\tau_{p}=0\}$ and $D_{\rho,r}\coloneqq\{\xi\in\mathcal{X}\colon\rho_{r}=0\}$. The jump set of $\mathcal{H}$ is
$D\coloneqq\big(\cup_{p\in\mathcal{V}}D_{\tau,p}\big)\cup\big(\cup_{r\in[M^{\ast}]}D_{\rho,r}\big).$ (26)
The difference inclusion governing the jumps of $\mathcal{H}$ is $\xi^{+}\in G(\xi)$, where the set-valued jump map $G\colon\mathcal{X}\rightrightarrows\mathcal{X}$ is
$G(\xi)$ $\coloneqq\{G_{\tau,p}(\xi)\colon\xi\in D_{\tau,p}\text{ for some }p\in\mathcal{V}\}\cup\{G_{\rho,r}(\xi)\colon\xi\in D_{\rho,r}\text{ for some }r\in[M^{\ast}]\},$ (27)
$G_{\tau,p}(\xi)$ $\coloneqq\left[\begin{array}[]{c}x^{\circ}\\ \left[\tilde{\eta}_{1}^{\top},...,\tilde{\eta}_{p-1}^{\top},0_{m}^{\top},\tilde{\eta}_{p+1}^{\top},...,\tilde{\eta}_{N}^{\top}\right]^{\top}\\ \tilde{\zeta}\\ \left[\tau_{1},...,\tau_{p-1},[T_{1}^{p},T_{2}^{p}],\tau_{p+1},...,\tau_{N}\right]^{\top}\\ \rho\end{array}\right],\quad G_{\rho,r}(\xi)\coloneqq\left[\begin{array}[]{c}x^{\circ}\\ \tilde{\eta}\\ \left[\tilde{\zeta}_{1}^{\top},...,\tilde{\zeta}_{r-1}^{\top},0_{mN}^{\top},\tilde{\zeta}_{r+1}^{\top},...,\tilde{\zeta}_{M^{\ast}}^{\top}\right]^{\top}\\ \tau\\ \left[\rho_{1},...,\rho_{r-1},[T_{3}^{r},T_{4}^{r}],\rho_{r+1},...,\rho_{M^{\ast}}\right]^{\top}\end{array}\right].$
@@
$\mathcal{A}\coloneqq\big\{\xi\in\mathcal{X}\colon(x^{\circ},\tilde{\eta},\tilde{\zeta})=(0_{n(N-1)},0_{mN},0_{mNM^{\ast}})\big\}$ (28)
@@
Under the construction of $\mathcal{H}$, the flow set $C$ and jump set $D$ are closed. Additionally, the flow map $f$ is continuous. The jump map $G$ is outer semi-continuous and locally bounded. As a result, $\mathcal{H}$ satisfies the hybrid basic conditions [Goebel.Sanfelice.ea2012, Assumption 6.5], and [Goebel.Sanfelice.ea2012, Theorem 6.8] implies $\mathcal{H}$ is nominally well-posed.
5 Stability Analysis